# T2 of Thoughts: Temperature Tree Elicits Reasoning in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools in artificial intelligence, especially in complex decision-making scenarios, but their static problem-solving strategies often limit their adaptability to dynamic environments. We explore the enhancement of reasoning capabilities in LLMs through Temperature Tree (T2) prompting via Particle Swarm Optimization, termed as T2 of Thoughts (T2oT). The primary focus is on enhancing decision-making processes by dynamically adjusting search parameters, especially temperature, to improve accuracy without increasing computational demands. We empirically validate that our hybrid T2oT approach yields enhancements in, single-solution accuracy, multi-solution generation and text generation quality. Our findings suggest that while dynamic search depth adjustments based on temperature can yield mixed results, a fixed search depth, when coupled with T2oT’s adaptive capabilities, provides a more reliable and versatile problem-solving strategy. This work highlights the potential for future explorations in optimizing algorithmic interactions with foundational language models, particularly illustrated by our development for the Game of 24 and Creative Writing tasks.
1 Introduction
Large Language models (LLMs) are increasingly employed across a broad spectrum of Natural Language Processing (NLP) tasks, including machine translation [1], summarization [2], and question answering [3]. Recent years a number of language modes such as GPT [4, 5, 6, 7], LLaMa [8, 9] are developed in a rapid speed. However, they are typically constrained to sequential, token-by-token processing during reasoning. These methods, while efficient, may be insufficient in scenarios that require forward-thinking or strategic planning, where early decisions significantly impact results.
@@ 1 Introduction
However, in these prompting methods, a significant parameter, temperature is often set as fixed [14, 15], which may introduce a number of limitations. To control the model generation, temperature sampling [16] is used to control the decoding process and influence the model performance by adjusting the probability distribution of the next token to be generated. To achieve a balance between generation quality and diversity, temperature sampling should be used properly [17]. In LLMs, the temperature primarily controls the randomness of text generation. A lower temperature value makes the model more likely to choose the conservative paths, leading to more deterministic and consistent outputs. Conversely, a higher temperature value increases the likelihood of selecting aggressive paths, resulting in more diverse and unpredictable text. As a result, fixed temperature restricts the flexibility to modulate the randomness and diversity of the responses based on different tasks. It has been demonstrated that using a fixed temperature is not the optimal choice in many cases, regardless of the type of language tasks being executed by the model [18].
@@ 1.1 Related Work
2 Methodology
2.1 Problem Formulation
@@
To address this, we propose T2oT for problem solving. The primary objective is to dynamically adjust the temperature parameter, thereby modulating the randomness and diversity of the responses based on the reasoning state. This dynamic adjustment is governed by PSO, which leverages personal best and global best values to fine-tune the temperature. Practically, T2oT can derive multiple trees. For each tree, it adjusts the temperature at each step of the reasoning process based on its own best evaluation value (personal best) and the best evaluation value among all the trees (global best). The problem can be formally defined as follows:
@@
$T_{i}[n]=w_{0}\cdot T_{i}[n-1]+\alpha_{1}\cdot(pb_{i}[n-1]-x_{i}[n])+\alpha_{2}\cdot(gb[n-1]-x_{i}[n])$ (1)
where:
  • $n$ is the n-th reasoning step.
  • $T$ is the temperature.
  • $i$ is the index of the current tree
  • $pb$ is the personal best value.
  • $gb$ is the global best value.
  • $x$ is the evaluation value.
  • $w_{0}$ is the inertial weight.
  • $\alpha_{1}$ and $\alpha_{2}$ are the acceleration coefficients for personal and global bests, respectively.
@@
As illustrated in Algorithm 1, we present the T2oT framework, where the temperature scale is dynamically adjusted. Throughout the reasoning process, the temperature undergoes continuous iteration in each tree.
@@
2.2 Theoretical Analysis
In this section, we have meticulously delineated the mathematical underpinnings of the T2oT algorithm, providing an analysis of its expectations, convergence, error bounds, and stability. This rigorous examination not only underscores the robustness and reliability of T2oT, but also paves the way for further exploration and optimization in large language model applications. More details refer to SupplementA.1.
Expectation and Variance in T2
To derive the expectations of the personal best ($pb$) and global best ($gb$).
The personal best for a particle $n$ at iteration $t+1$ is updated as:
$pb_{n}[t+1]=\begin{cases}x_{n}[t+1],&\text{if }f(x_{n}[t+1])<f(pb_{n}[t])\\ pb_{n}[t],&\text{otherwise}\end{cases}$ (2)
The expected value of the personal best can be written as:
$\mathbb{E}[pb_{n}[t+1]]=\mathbb{E}[\min(pb_{n}[t],x_{n}[t+1])]$ (3)
Using the law of total expectation, we have:
$\mathbb{E}[pb_{n}[t+1]]=\mathbb{E}[pb_{n}[t]]\cdot\Pr(x_{n}[t+1]\geq pb_{n}[t])+\mathbb{E}[x_{n}[t+1]]\cdot\Pr(x_{n}[t+1]<pb_{n}[t])$ (4)
Assuming a uniform distribution for the positions, the probabilities can be simplified. Let $U$ be the uniform distribution over the search space:
$\Pr(x_{n}[t+1]<pb_{n}[t])=F(pb_{n}[t])$ (5)
where $F$ is the cumulative distribution function of the uniform distribution.
The expected value of $x_{n}(t+1)$ is:
[…]
Thus, the expected value of the personal best becomes:
$\mathbb{E}[pb_{n}[t+1]]=\mathbb{E}[pb_{n}[t]]\cdot(1-F(pb_{n}[t]))+\frac{a+b}{2}\cdot F(pb_{n}[t])$ (7)
Expectation of Global Best (gb): The global best at iteration $t+1$ is updated as:
$gb[t+1]=\min(gb[t],\min_{n}f(x_{n}[t+1]))$ (8)
Convergence Analysis: Assuming that the algorithm converges to a stable temperature $T^{*}$, then we can get:
$T^{*}=\frac{\alpha_{1}\cdot(pb^{*}-x^{*})+\alpha_{2}\cdot(gb^{*}-x^{*})}{1-w_{0}}$ (9)
Error Bounds: Let $e_{n}(t)$ denote the error in temperature adjustment for particle $n$ at iteration $t$. The error can be defined as:
$e_{n}[t]=T_{n}[t]-T^{*}$ (10)
$E_{\max}=\frac{\alpha_{1}\cdot\max|pb_{n}[t]-x_{n}|+\alpha_{2}\cdot\max|gb[t]-x_{n}|}{1-w_{0}}$ (11)
This error bound provides an upper limit on the deviation of the temperature from the optimal value $T^{*}$, $\alpha_{1},\alpha_{2}$, and $w_{o}$ are hyperparameters. For more details refer to A.1.2.
Assumption 1.
The T2oT algorithm is applied to a dataset of reasoning steps $(x_{i},t_{i})$, where $x_{i}$ represents the input at step $i$, and $t_{i}$ is the corresponding temperature. The pairs $(x_{i},t_{i})$ are drawn i.i.d. from an unknown distribution $\mathcal{D}$.
Assumption 2.
The temperature adjustment loss $L_{temp}$ is $L$-Lipschitz continuous with respect to the temperature $T$.
Stability and Robustness: The stability of the T2oT algorithm is ensured if the inertial weight $w_{0}$, and the acceleration coefficients $\alpha_{1}$ and $\alpha_{2}$ are chosen such that $w_{0}<1$.
3 Experiments
3.1 Game of 24
3.1.1 Experimental Setup
@@
To evaluate our method, we employed the same task as used in ToT: Game of 24 in GPT-4. This game involves using arithmetic operations to manipulate four numbers to arrive at the number 24, a common benchmark in studies of computational problem solving and algorithmic efficiency. This specific task selection helps in directly comparing the performance and efficacy of T2oT against established results in ToT, providing a clear metric for improvement or competitiveness in problem-solving scenarios. To evaluate the enhancements of T2oT, we conducted two distinct types of experiments. The first experiment is designed to evaluate the performance of our method across 50 games on each method, comparing the accuracy of T2oT against ToT. The second experiment focused specifically on T2oT’s capability to explore multiple solutions. We selected a specific instance of a game "7, 5, 2, 6" which have multiple solutions and run both methods on this instance 10 times each and track the frequency of distinct solutions produced by each run.
@@
In order to set up a reliable experiment plan for T2oT, to test its performance when applied into a large language model, a newly generated dataset is necessary. To align with our experiments, the dataset for this task was algorithmically generated. Each instance of the Game of 24 consists of four numbers. For more details refer to A.2.2.
@@
For our experiments, we use ToT as the baseline and the same setup in our T2oT. T2oT breaks down the thought process into three steps. We set the depth of search to 3. Utilizing a breadth-first search (BFS) methodology within the framework, we consistently select and advance the top five most promising solutions at every stage of the decision-making process. As shown in Figure 2, this selection is evaluated by the language model prompt, which classifies each potential solution into categories: "sure", "maybe" or "impossible" based on the likelihood of reaching the total of 24. The primary objective remains to progress viable partial solutions that can be substantiated with a limited number of predictive steps, eliminate those solutions that are evidently unfeasible through practical judgments such as being "too large" or "too small," and continue to explore those deemed as "maybe". We perform value sampling three times for each conceptual step in the thought process.
@@
3.1.2 Results
Table 1: Game of 24 Success Rates
@@ 3.1.2 Results
Table 2: Game of 24 Multiple Solutions Results
@@
As shown in Table 2, the success rate for T2oT is 80%, which is higher than ToT at 72%. This indicates an improvement of 8% increase in success. T2oT appears to provide a more effective search strategy than ToT. The increased accuracy suggests that T2oT may be better at navigating the problem space or more efficient at finding solutions that meet the goal. This could be due to better exploration of the search space or more effective pruning of less promising paths.
The result of the second experiment is shown in Table 2. ToT is able to generate two distinct solution types for the given game, with the primary solution type occurring with a high frequency of 0.9 and a secondary, less frequent type at 0.1. In contrast, T2oT demonstrates a superior ability to diversify its approach, producing three different types of solutions with a more balanced distribution of frequencies at 0.5, 0.3, and 0.2, respectively. This indicates that T2oT not only increases the number of potential solutions explored but also promotes a more equitable exploration across these solutions, preventing over-concentration on a dominant solution path. More details about this experiment can refer to A.2.3.
3.2 Creative Writing
3.2.1 Experimental Setup
@@
Figure 3: T2oT in Creative Writing. After generating five plans by prompting, dynamically adjust the temperature for producing passages based on the highest score. Full passage can be seen in A.2.5
@@
For our experiments, we also utilize ToT as the baseline, employing a comparable setup in our T2oT. As shown in Figure 3, T2oT breaks down the thought process into two steps (depth=2): Firstly, the language model generates 5 different plans, then conducts 5 rounds of voting to determine the best plan. Secondly, the majority choice is used to write the output passage, following the same generate-and-vote procedure. As we find that GPT can follow the input constraints most of the time, and since the original evaluation method can lead to inflated scores in the first step of plan assessment, and the 10-point scale results in many similar scores, making it difficult to distinguish among plans, we use a GPT zero-shot prompt to provide a 1-100 scalar score: "Analyze the following passage in detail. Consider the clarity, structure, argument coherence, and style of the writing. (……) Conclude with: ’Thus, the coherency score is s’, where s is an integer from 0 to 100, aiming for a normal distribution of scores with an average of 50." (see full prompt inA.2.4). This evaluation method aims for a normal distribution of scores with an average score of 50.
@@
3.2.2 Results
Figure 4: Creative Writing results.
Figure 4(a) presents the average GPT-4 coherency scores across various experimental setups, illustrating that T2oT, with an average score of 71.4, produces more coherent passages compared to IO (62.28), CoT (64.8) and ToT (67.5). The T2oT appears to provide a more effective search strategy and offer a good dynamic balance: lowering the temperature to achieve higher textual consistency when plan consistency is low, and raising the temperature to enhance creativity when plan consistency is high. Figure 4(b) confirms the trend by showing that humans prefer ToT over T²oT in 9 out of 50 passage pairs, while preferring T²oT over ToT in 20 pairs, with the remaining 21 pairs judged as similarly coherent.
4 Discussion
@@
Limitations
While our method can dynamically adjust the temperature and demonstrate significant efficiency improvements, our algorithm is not a gradient-based and learnable algorithm. In our task, the inertial weight, acceleration coefficient for personal best, and acceleration coefficient for global best in the T2oT are all fixed values. This means they may need manual adjustment depending on the specific task. Therefore, integrating neural networks with the reasoning process can allow for learning these parameter selections based on performance outcomes.
@@
In this paper, we presents T2oT, a novel prompting method that dynamically adjusts the temperature parameter during inference in large language models, significantly enhancing both accuracy and diversity of solutions generated by GPT-4. Through comprehensive empirical validation, T2oT demonstrates superior performance over the static temperature method ToT, achieving higher single-solution accuracy and improved multi-solution generation in tasks such as the Game of 24, as well as better coherency in Creative Writing. This integration of heuristic algorithms with artificial intelligence exemplifies a promising approach for developing more adaptive and efficient language model prompting techniques.
Appendix A Appendix
A.1 Explanation for T2oT
Expected Value of Global Best
The global best at iteration $t+1$ is updated as:
$gb[t+1]=\min(gb[t],\min_{n}f(x_{n}[t+1]))$ (12)
The expected value of the global best can be written as:
$\mathbb{E}[gb[t+1]]=\mathbb{E}[\min(gb[t],\min_{n}x_{n}[t+1])]$ (13)
Using the law of total expectation, we have:
$\mathbb{E}[gb[t+1]]=\mathbb{E}[gb[t]]\cdot\Pr(\min_{n}x_{n}[t+1]\geq gb[t])+\mathbb{E}[\min_{n}x_{n}[t+1]]\cdot\Pr(\min_{n}x_{n}[t+1]<gb[t])$ (14)
For the global best:
$\Pr(\min_{n}x_{n}[t+1]<gb[t])=1-(1-F(gb[t]))^{N}$ (15)
where $N$ is the number of trees.
Putting these together:
$\mathbb{E}[gb[t+1]]=\mathbb{E}[gb[t]]\cdot\left(1-(1-F(gb[t]))^{N}\right)+\mathbb{E}[\min_{n}x_{n}[t+1]]\cdot(1-F(gb[t])^{N})$ (16)
A.1.1 Convergence Analysis
The convergence of the T2 algorithm can be analyzed by examining the behavior of the particles as $t\to\infty$. Assuming that the algorithm converges to a stable temperature $T^{*}$, then we can get:
$T^{*}=w_{0}\cdot T^{*}+\alpha_{1}\cdot(pb^{*}-x^{*})+\alpha_{2}\cdot(gb^{*}-x^{*})$ (17)
Solving for $T^{*}$ gives:
$T^{*}=\frac{\alpha_{1}\cdot(pb^{*}-x^{*})+\alpha_{2}\cdot(gb^{*}-x^{*})}{1-w_{0}}$ (18)
This equation indicates the conditions under which the temperature converges to a stable value.
A.1.2 Error Bounds
To ensure the reliability of the T2 algorithm, it is crucial to establish error bounds for the temperature adjustments. Let $e_{n}(t)$ denote the error in temperature adjustment for tree $n$ at iteration $t$. The error can be defined as:
$e_{n}[t]=T_{n}[t]-T^{*}$ (19)
The error bounds can then be derived by analyzing the propagation of errors over iterations:
$|e_{n}[t+1]|\leq w_{0}\cdot|e_{n}[t]|+\alpha_{1}\cdot|pb_{n}-x_{n}-T^{*}|+\alpha_{2}\cdot|gb-x_{n}-T^{*}|$ (20)
Given that $|pb_{n}-x_{n}|$ and $|gb-x_{n}|$ are bounded, we can define a maximum error bound $E_{\max}$:
$E_{\max}=\frac{\alpha_{1}\cdot\max|pb_{n}[t]-x_{n}|+\alpha_{2}\cdot\max|gb[t]-x_{n}|}{1-w_{0}}$ (21)
This error bound provides an upper limit on the deviation of the temperature from the optimal value $T^{*}$.
A.1.3 Stability and Robustness
The stability of the T2 algorithm is ensured if the inertial weight $w_{0}$, and the acceleration coefficients $\alpha_{1}$ and $\alpha_{2}$ are chosen such that $w_{0}<1$. This ensures that the error diminishes over iterations, leading to convergence.
Robustness can be analyzed by introducing perturbations or noise into the temperature adjustments and studying their impact on the convergence properties. Let $\eta(t)$ represent a noise term:
$T_{n}[t+1]=w_{0}\cdot T_{n}[t]+\alpha_{1}\cdot(pb_{n}[t]-x_{n})+\alpha_{2}\cdot(gb[t]-x_{n})+\eta[t]$ (22)
To ensure robustness, the magnitude of $\eta(t)$ should be controlled such that it does not disrupt the convergence process.
A.2 Experiments
A.2.1 One-Tree T2oT
Figure 5: The reasoning process of T2oT when the number of trees is set to one.
Figure 5 illustrates the reasoning process of T2oT when the number of trees is limited to one, showing how the initial temperature adjusts dynamically through multiple steps based on the quality of thoughts—categorized as good, average, and bad—to produce the final output. Specifically, in each reasoning step, the adjusted temperature is iterated according to algorithm 1.
A.2.2 Dataset Generation of Game of 24
Algorithm 2 Generating Game of 24
1:A set of four numbers: game_set, final result is 24
2:for every first and second elements in the input array, apply +, -, $\times$, and $\div$ to the first two elements and calculate the remaining.
3:for the remaining and the other elements in the array, merge them to build a new array: new_set
4:if remaining $\neq$ 24 then
5:     execute the Algorithm of generating Game of 24 based on the new_set
6:else if remaining $=$ 24 then
7:     return the original game_set
8:end if
Algorithm 2 is used to generate our dataset for the experiment in Game of 24.
A.2.3 Details about Multiple Solution of Game of 24
Table 3: Details of Game of 24 Multiple Solutions
Method Solution Types Frequency
ToT (7 - 5) * 2 * 6 = 24 0.9
ToT 5 + 7 + ( 2 * 6 ) = 24 0.1
ToT (7 - 5 + 2) * 6 = 24 0
T2oT (ours) (7 - 5) * 2 * 6 = 24 0.3
T2oT (ours) 5 + 7 + ( 2 * 6 ) = 24 0.5
T2oT (ours) (7 - 5 + 2) * 6 = 24 0.2
Table 3 contains the specific solutions of the game "7, 5, 2, 6" and their frequency in our multiple solution experiments.
A.2.4 Prompt Used for Creative Writing Scores
prompt = "Analyze the following passage in detail. Consider the clarity, structure, argument coherence, and style of the writing. Use the following scale to assign a coherency score:
- 0-10: The text is mostly incoherent or very poorly structured. Major issues pervade the text, making it nearly impossible to understand.
- 11-30: The text has significant issues in clarity or structure that heavily impede understanding. Many parts of the text do not flow logically.
- 31-50: The text is adequately coherent and structured but lacks refinement and contains noticeable errors that affect readability.
- 51-70: The text is coherent and well-structured with some room for improvement. Minor errors or stylistic inconsistencies are present.
- 71-90: The text is very well-structured and coherent. It is clear, with only slight imperfections in style or structure.
- 91-100: The text demonstrates exceptional clarity and structure, showcasing excellent coherence and nearly flawless execution.
Conclude with: "Thus, the coherency score is s", where s is an integer from 0 to 100, aiming for a normal distribution of scores with an average of 50."
A.2.5 Creative Writing Passage Example
Jake was generally a careful man, one who adhered to his routine and went about his days with precision. He woke up at the same time each day, enjoyed a cup of coffee in the morning, and left for work punctually. However, this morning was not a typical one. Today arrived with a crash of my car through the garage door.
The car accident was just the first in a series of unfortunate events Jake was to experience that day. At work, he was tasked with predicting the outcome of a significant business deal. The pressure was immense and the task, unpredictable. It was akin to standing over a pot of kernels waiting for them to pop - a daunting spectacle of uncertainty. It’s never comforting to know that your fate depends on something as unpredictable as the popping of corn.
After a gruelling day, Jake decided to unwind at the beach. The beach had always been his sanctuary, a place where he could relax and enjoy some fresh air. But today, it felt different. The sand was irritating him, getting into his shoes and his clothes, and the sun was blazing, causing him discomfort. He was disappointed when he found the beach to be so sandy and the sun so sunny.
As Jake walked back home, he reflected on his day. He had always relied on his meticulous planning and careful approach to navigate through life. But today’s events made him realize that life is unpredictable and sometimes, it required more than just careful planning. It required courage to face the unknown and a touch of stupidity to venture into it. Courage and stupidity were all he had.

### v2

@@ Abstract
Large Language Models (LLMs) have emerged as powerful tools in artificial intelligence, especially in complex decision-making scenarios, but their static problem-solving strategies often limit their adaptability to dynamic environments. We explore the enhancement of reasoning capabilities in LLMs through Temperature Tree (T2) prompting via a heuristic algorithm, termed as T2 of Thoughts (T2oT). The primary focus is on enhancing decision-making processes by dynamically adjusting search parameters, especially temperature, to improve accuracy without increasing computational demands. We empirically validate that our hybrid T2oT approach yields enhancements in, single-solution accuracy, multi-solution generation and text generation quality. Our findings suggest that while dynamic search depth adjustments based on temperature adaption can yield mixed results, a fixed search depth, when coupled with T2oT’s adaptive capabilities, provides a more reliable and versatile problem-solving strategy. This work highlights the potential for future explorations in optimizing algorithmic interactions with foundational language models, particularly illustrated by our development for the Game of 24 and Creative Writing tasks.
Index Terms:
large language model, prompt method, heuristic optimization
I Introduction
Large Language models (LLMs) are increasingly employed across a broad spectrum of Natural Language Processing (NLP) tasks, including machine translation [1], summarization [2], and question answering [3]. In Recent years, a number of language models such as GPT [4, 5, 6, 7], LLaMa [8, 9] are developed in a rapid speed. However, they are typically constrained to sequential, token-by-token processing during reasoning. These methods, while efficient, may be insufficient in scenarios that require forward-thinking or strategic planning, where early decisions significantly impact results.
@@ I Introduction
However, in these prompting methods, a significant parameter, temperature, is often set as fixed [14, 15], which may introduce a number of limitations. To control the model generation, temperature sampling [16] is used to control the decoding process and influence the model performance by adjusting the probability distribution of the next token to be generated. To achieve a balance between generation quality and diversity, temperature sampling should be used properly [17]. In LLMs, the temperature primarily controls the randomness of text generation. A lower temperature value makes the model more likely to choose the conservative paths, leading to more deterministic and consistent outputs. Conversely, a higher temperature value increases the likelihood of selecting aggressive paths, resulting in more diverse and unpredictable text. As a result, fixed temperature restricts the flexibility to modulate the randomness and diversity of the responses based on different tasks. It has been demonstrated that using a fixed temperature is not the optimal choice in many cases, regardless of the type of language tasks being executed by the model [18].
@@ II Related Work
III Methodology
Problem Formulation.
@@ III Methodology
To address this, we propose T2oT for problem solving in Figure 1. The primary objective is to dynamically adjust the temperature parameter, thereby modulating the randomness and diversity of the responses based on the reasoning state. This dynamic adjustment is governed by PSO, which leverages personal best and global best values to fine-tune the temperature. Practically, T2oT can derive multiple trees. For each tree, it adjusts the temperature at each step of the reasoning process based on its own best evaluation value (personal best) and the best evaluation value among all the trees (global best). The problem can be formally defined as follows:
@@
$T_{i}[n]$ $=w_{0}\cdot T_{i}[n-1]$ (1)
$+\alpha_{1}\cdot(pb_{i}[n-1]-x_{i}[n])$
$+\alpha_{2}\cdot(gb[n-1]-x_{i}[n])$
In the Equation 1, $n$ represents the n-th reasoning step, $T$ denotes the temperature, $i$ is the index of the current tree, $pb$ stands for the personal best value, $gb$ represents the global best value, $x$ indicates the evaluation value, $w_{0}$ is the inertial weight, and $\alpha_{1}$ and $\alpha_{2}$ are the acceleration coefficients for personal and global bests, respectively.
@@
The T2oT framework efficiently solves the temperature optimization problem by dynamically adjusting the temperature scale at each iteration, as described in Algorithm 1. This iterative process ensures that the pipeline converges to the correct temperature $T^{*}$ while avoiding significant errors. Through careful tuning of hyperparameters $\alpha_{1}$, $\alpha_{2}$, and $w_{0}$, the framework balances local and global factors, enabling accurate and efficient convergence to the optimal solution with minimal error.
IV Experiment
IV-A Game of 24
IV-A1 Experimental Setup
@@
To evaluate our method, we employed the same task as used in ToT: Game of 24 in GPT-4. This game involves using arithmetic operations to manipulate four numbers to arrive at the number 24, a common benchmark in studies of computational problem solving and algorithmic efficiency. This specific task selection helps in directly comparing the performance and efficacy of T2oT against established results in ToT, providing a clear metric for improvement or competitiveness in problem-solving scenarios. To evaluate the enhancements of T2oT, we conducted two distinct types of experiments. The first experiment is designed to evaluate the performance of our method across 50 games on each method, comparing the accuracy of T2oT against ToT. The second experiment focused specifically on T2oT’s capability to explore multiple solutions. We selected a specific instance of a game ”7, 5, 2, 6” which have multiple solutions and run both methods on this instance 10 times each and track the frequency of distinct solutions produced by each run.
@@
In order to set up a reliable experiment plan for T2oT, to test its performance when applied into a large language model, a newly generated dataset is necessary. To align with our experiments, the dataset for this task was algorithmically generated. Each instance of the Game of 24 consists of four numbers.
@@
For our experiments, we use ToT as the baseline and the same setup in our T2oT. T2oT breaks down the thought process into three steps. We set the depth of search to 3. Utilizing a breadth-first search (BFS) methodology within the framework, we consistently select and advance the top five most promising solutions at every stage of the decision-making process. As shown in Figure 2, this selection is evaluated by the language model prompt, which classifies each potential solution into categories: ”sure”, ”maybe” or ”impossible” based on the likelihood of reaching the total of 24. The primary objective remains to progress viable partial solutions that can be substantiated with a limited number of predictive steps, eliminate those solutions that are evidently unfeasible through practical judgments such as being ”too large” or ”too small,” and continue to explore those deemed as ”maybe”. We perform value sampling three times for each conceptual step in the thought process.
Figure 3: T2oT in Creative Writing. After generating five plans by prompting, dynamically adjust the temperature for producing passages based on the highest score.
@@
IV-A2 Results
TABLE I: Game of 24 Success Rates
@@ IV-A2 Results
TABLE II: Game of 24 Multiple Solutions Results
@@
As shown in Table I, IO and CoT prompt methods perform badly on this task. The success rate for T2oT is 80%, which is higher than ToT at 72%. This indicates an improvement of 8% increase in success. T2oT appears to provide a more effective search strategy than ToT. The increased accuracy suggests that T2oT may be better at navigating the problem space or more efficient at finding solutions that meet the goal. This could be due to better exploration of the search space or more effective pruning of less promising paths.
The result of the second experiment is shown in Table II. ToT is able to generate two distinct solution types for the given game, with the primary solution type occurring with a high frequency of 0.9 and a secondary, less frequent type at 0.1. In contrast, T2oT demonstrates a superior ability to diversify its approach, producing three different types of solutions with a more balanced distribution of frequencies at 0.5, 0.3, and 0.2, respectively. This indicates that T2oT not only increases the number of potential solutions explored but also promotes a more equitable exploration across these solutions, preventing over-concentration on a dominant solution path.
IV-B Creative Writing
IV-B1 Experimental Setup
Figure 4: Creative Writing results.
@@ IV-B1 Experimental Setup
For our experiments, we also utilize ToT as the baseline, employing a comparable setup in our T2oT. As shown in Figure 3, T2oT breaks down the thought process into two steps (depth = 2): Firstly, the language model generates 5 different plans, then conducts 5 rounds of voting to determine the best plan. Secondly, the majority choice is used to write the output passage, following the same generate-and-vote procedure. As we find that GPT can follow the input constraints most of the time, and since the original evaluation method can lead to inflated scores in the first step of plan assessment, and the 10-point scale results in many similar scores, making it difficult to distinguish among plans, we use a GPT zero-shot prompt to provide a 1-100 scalar score: ”Analyze the following passage in detail. Consider the clarity, structure, argument coherence, and style of the writing. (……) Conclude with: ’Thus, the coherency score is s’, where s is an integer from 0 to 100, aiming for a normal distribution of scores with an average of 50.” This evaluation method aims for a normal distribution of scores with an average score of 50.
@@
IV-B2 Results
Figure 4(a) presents the average GPT-4 coherency scores across various experimental setups in GPT-4-0613 , illustrating that T2oT, with an average score of 71.4, produces more coherent passages compared to IO (62.28), CoT (64.8) and ToT (67.5). While Figure 4(b) illustrates that in Claude-3.5-haiku-20241022, our T2oT has an average score of 68.3, which is also higher than IO (58.7), CoT (59.4) and ToT (63.7). It is worth noting that in the Creative Writing experiments conducted with Claude in Figure 4(b), we still used GPT-4-0613 to evaluate the results to ensure consistency with the evaluation metric used in experiments conducted with GPT-4 in Figure 4(a). The T2oT appears to provide a more effective search strategy and offer a good dynamic balance: lowering the temperature to achieve higher textual consistency when plan consistency is low, and raising the temperature to enhance creativity when plan consistency is high. Figure 4(c) confirms the trend by showing that in the experimets of GPT-4, humans prefer the results generated by ToT over T²oT in 9 out of 50 passage pairs, while preferring T²oT over ToT in 20 pairs, with the remaining 21 pairs judged as similarly coherent. Figure 4(d) shows that humans preferred the results generated by ToT in 7 pairs, while preferring T²oT in 24 pairs, with the remaining 19 pairs judged as similarly coherent.
TABLE III: Evaluation with non-default temperature ToT
Method Score (Mean ± Std)
IO 61.41 ± 9.48
CoT 63.71 ± 11.00
ToT (random temperature) 65.26 ± 10.36
ToT 67.57 ± 10.31
T²oT 71.59 ± 8.79
Table III presents the results of using ToT with randomly sampled temperatures from the range (0,1) at each inference step, compared to default temperature ToT and our T²oT method. Since our approach dynamically adjusts the temperature during inference, this experiment demonstrates that random temperature adjustment does not lead to performance improvements. In contrast, the temperature adjustments made by our algorithm contribute to enhanced performance.
V Cost and efficiency
Under the single-particle conditions used in this paper, the computation cost of using T2oT is very similar to that of ToT. Table IV shows that in Game of 24, solving a problem with T2oT and ToT both requires about 5.5k completion tokens, which is very similar to the tokens required by 100 CoT trials and selecting the best result. This shows that the number of tokens required to scoring the results in T2oT is very small compared to the computation.
TABLE IV: Game of 24 Cost
Method Generate/Prompt tokens Cost per case
IO 1.8k / 1.0k 0.13
CoT 6.7k / 2.2k 0.47
ToT 5.5k / 1.4k 0.74
T2oT(ours) 5.5k / 1.6k 0.80
In Creative Writing (Table V below), the number of completion tokens required by T2oT and ToT are also very similar. As a result, in our experiments, our single-particle T2oT has a significant performance improvement than ToT without increasing the number of completion tokens and with less increase in cost.
TABLE V: Creative Writing Cost
Method Generate/Prompt tokens Cost per case
IO 0.9k / 0.4k 0.13
CoT 0.9k / 0.4k 0.47
ToT 4k / 2.9k 0.74
T2oT(ours) 4k / 3.8k 1.01
VI Discussion
@@ VII Conclusion
In this paper, we present T2oT, a novel reasoning method that dynamically adjusts the temperature parameter during inference in large language models, significantly enhancing both the accuracy and diversity of solutions generated by GPT-4. Through comprehensive empirical validation, T2oT demonstrates superior performance over the static temperature method ToT, achieving higher single-solution accuracy and improved multi-solution generation in tasks such as the Game of 24, as well as better coherency in Creative Writing. This integration of heuristic algorithms with artificial intelligence exemplifies a promising approach for developing more adaptive and efficient language model prompting techniques.
VIII Limitations
While the proposed T2oT algorithm demonstrates significant improvements in reasoning and solution diversity compared to traditional fixed-temperature methods, it still has several limitations. First, the manual tuning of hyperparameters such as inertial weight and acceleration coefficients can be suboptimal in different tasks. These fixed parameters may limit the adaptability of the algorithm in dynamic environments where real-time adjustments could lead to better performance.
Another limitation is the absence of a learnable mechanism for adjusting the temperature parameter. The current approach relies on a heuristic strategy based on particle swarm optimization (PSO), which, although effective, may not always provide the most efficient temperature adjustments across all problem domains. Future work could explore integrating reinforcement learning techniques or neural networks to automatically learn these hyperparameters based on feedback from the model’s performance with multiple trees.
Furthermore, while T2oT has demonstrated robust performance in structured tasks such as the Game of 24 and Creative Writing, its generalizability to more complex or less structured tasks, particularly those involving multimodal data, remains to be fully tested. The evaluation of its effectiveness in such domains would require further research.